# Clustering Malware at Scale: A First Full-Benchmark Study

### v1

@@ Abstract
Recent years have shown that malware attacks still happen with high frequency. Malware experts seek to categorize and classify incoming samples to confirm their trustworthiness or prove their maliciousness. One of the ways in which groups of malware samples can be identified is through malware clustering. Despite the efforts of the community, malware clustering which incorporates benign samples has been under-explored. Moreover, despite the availability of larger public benchmark malware datasets, malware clustering studies have avoided fully utilizing these datasets in their experiments, often resorting to small datasets with only a few families. Additionally, the current state-of-the-art solutions for malware clustering remain unclear. In our study, we evaluate malware clustering quality and establish the state-of-the-art on Bodmas and Ember - two large public benchmark malware datasets. Ours is the first study of malware clustering performed on whole malware benchmark datasets. Additionally, we extend the malware clustering task by incorporating benign samples. Our results indicate that incorporating benign samples does not significantly degrade clustering quality. We find that there are significant differences in the quality of the created clusters between Ember and Bodmas, as well as a private industry dataset. Contrary to popular opinion, our top clustering performers are K-Means and BIRCH, with DBSCAN and HAC falling behind.
@@ 1 Introduction
  • Our experiments show that the malware clustering quality is highly dependent on the (sample) composition of the dataset - with the Ember dataset achieving sub-par results in Homogeneity compared to Bodmas. Our ablation study shows only a small Homogeneity increase when increasing the number of components, ruling out the number of clusters as a significant contributor to the low Ember cluster quality.
@@ 4.1 Quantitative Results
In the top-left graph of Figure 2, we can see that, on average, the clustering algorithms achieved the highest Homogeneity on Bodmas - above 70%. Security, the private real-world industry dataset, achieved second place with Homogeneity being, on average, around 40%. Surprisingly, Ember was the worst of the three datasets to cluster, with mean Homogeneity only above 20%.
Looking at maximum Homogeneity (in the top-right graph of Figure 2) achieved on the datasets, we can see that the difference between mean and maximum Homogeneity can often be quite large, above 20%. This is related to the relatively large spread of the number of clusters used in our experiments. Table 3 summarizes the best results achieved for the respective datasets. From it, we can see that K-Means consistently beat the other clustering algorithms in the race for best results. The best result overall, and the best result for Bodmas as well, was 92.5% Homogeneity achieved using UMAP representation. The best achieved Homogeneity on a real-world dataset, Security, was 77.64% - a relatively good result considering that this is the largest dataset in terms of samples and number of malware families. For Ember, all the clustering methods struggled to achieve a good clustering quality, with the best result being 42.66% Homogeneity. On all of the datasets, V-Measure stayed relatively low (8% to 16%) for the best results - a consequence of penalization through Completeness because of the utilization of more clusters than the reported number of malware families.
@@ 4.2 Ablation Experiment on the Number of Components
The results of the ablation experiment are reported in Table 4. For Bodmas, the Homogeneity seemed to increase with every increase in the number of components except for the case of Autoencoder and 30 components. For Ember, Homogeneity consistently increased with the increase in the number of components. The results for the Security dataset are the least consistent among the three datasets. In the case of PCA and 30 components, the Homogeneity decreased a little. On the other hand, for Autoencoder and 30 components, we saw an increase of more than 2% Homogeneity. For the case of 30 components, the Homogeneity increased in four out of six cases (of combinations of dataset and representation). The Homogeneity increased in all six cases for 50 components (compared to the 10-component baseline). Conclusions about this experiment will be drawn in Section 5.
@@ 5.1 Interpretation of Results
The results presented in Section 4 provide multiple insights. The results clearly show that Bodmas was the easiest dataset to cluster, achieving above 90% Homogeneity for the top result. Contrastingly, Ember was the dataset where the worst results in terms of Homogeneity were achieved (i.e., only reaching 42% for the best result on Ember). This constitutes a considerable variance in results. Both of these datasets are public benchmark datasets and share the same set of $2381$ Ember features. Moreover, they were gathered only roughly 2 years apart - from January 2017 to March 2018 for Ember and from August 2019 to September 2020 for Bodmas [yang2021bodmas]. Therefore, the main differentiating factor are the samples inside the datasets. Thus, our key finding is that malware clustering quality is highly dependent on the composition of the dataset.
Our work modified the malware clustering scenario by including benign samples. We can make conclusions based on the results observed in our experiments and in the experiments found in the existing work presented in Section 2.Although often performed on much smaller datasets with much fewer malware families, the existing work showed 98% Purity, Precision between 72% to 89%, 92% V-Measure, and Homogeneity of 73% and 79%. Thus, our observed results for Bodmas and Security align with the existing literature’s observations. We conclude that the inclusion of benign samples did not have a dramatic effect on the clustering quality in terms of Homogeneity. The result is different for Ember. As discussed earlier, the results on Ember are primarily affected by the composition of the dataset itself and therefore we see Ember as an exception to the conclusion.
@@
Third, one of our primary motivations behind the experiment was also to see the impact of the increase in the number of components on the Ember dataset. Unfortunately, even for Ember, for which we have seen the worst results in terms Homogeneity (in the main experiment), the increase in the number of components did not help much in increasing Homogeneity. Ideally, we would like to see that the number of components has a greater influence (i.e., at least 5%, ideally at least 10%) on the result Homogeneity, but this was not observed. This strengthens our previously stated argument that the influence of the samples (picked into the dataset) on the result Homogeneity cannot be overstated.
@@ 5.3 Future Work
A deeper investigation into the poor results on the Ember dataset is equally important. For instance, merging benign and malware samples from different datasets (e.g., Ember and Bodmas) and repeating the clustering experiments could offer better insights into malware clustering. This approach may help clarify the challenges associated with clustering different parts of these datasets.
@@ 6 Conclusion
We have established the state-of-the-art in malware clustering on large public malware benchmark datasets. The best results for Bodmas and Ember were achieved using K-Means and were 91% and 44% Homogeneity, respectively. Our results also indicate that K-Means and BIRCH achieve higher Homogeneity on these large datasets than DBSCAN and Hierarchical Agglomerative Clustering. Moreover, our findings show that the quality of malware clustering can vary considerably, even if they share the same set of features. This led us to conclude that malware clustering quality is highly dependent on the composition of the dataset. To examine the results more closely, we performed an ablation study, which increased the number of components of PCA and Autoencoder representations. The ablation results show that the increase in the number of components does not have a substantial effect on clustering Homogeneity, strengthening the result regarding dataset composition.

### v2

@@ Abstract
Recent years have shown that malware attacks still happen with high frequency. Malware experts seek to categorize and classify incoming samples to confirm their trustworthiness or prove their maliciousness. One of the ways in which groups of malware samples can be identified is through malware clustering. Despite the efforts of the community, malware clustering which incorporates benign samples has been under-explored. Moreover, despite the availability of larger public benchmark malware datasets, malware clustering studies have avoided fully utilizing these datasets in their experiments, often resorting to small datasets with only a few families. Additionally, the current state-of-the-art solutions for malware clustering remain unclear. In our study, we evaluate malware clustering quality and establish the state-of-the-art on Bodmas and Ember - two large public benchmark malware datasets. Ours is the first study of malware clustering performed on whole malware benchmark datasets. Additionally, we extend the malware clustering task by incorporating benign samples. Our results indicate that incorporating benign samples does not significantly degrade clustering quality. We find that there are differences in the quality of the created clusters between Ember and Bodmas, as well as a private industry dataset. Contrary to popular opinion, our top clustering performers are K-Means and BIRCH, with DBSCAN and HAC falling behind.
@@ 1 Introduction
  • Our experiments show that the malware clustering quality is dependent on the (sample) composition of the dataset - with the Ember dataset achieving sub-par results in Homogeneity compared to Bodmas. Our ablation study shows only a small Homogeneity increase when increasing the number of components, ruling out the number of clusters as a significant contributor to the low Ember cluster quality.
@@ 4.1 Quantitative Results
In the top-left graph of Figure 2, we can see that, on average, the clustering algorithms achieved the highest Homogeneity on Bodmas - above 70%. On Ember, the larger public benchmark malware dataset of the two, the clustering algorithms managed to achieve around 50% Homogeneity on average. Security, the private real-world industry dataset, achieved a solid third place with Homogeneity being, on average, around 40%.
Looking at maximum Homogeneity (in the top-right graph of Figure 2) achieved on the datasets, we can see that the difference between mean and maximum Homogeneity can often be quite large, above 20%. This is related to the relatively large spread of the number of clusters used in our experiments. Table 3 summarizes the best results achieved for the respective datasets. From it, we can see that K-Means consistently beat the other clustering algorithms in the race for best results. The best result overall, and the best result for Bodmas as well, was 92.5% Homogeneity achieved using UMAP representation. The best achieved Homogeneity on a real-world dataset, Security, was 77.64% - a relatively good result considering that this is the largest dataset in terms of samples and number of malware families. With regards to Ember, the clustering methods struggled to achieve a clustering quality that would be on par with Bodmas. Nevertheless, the best Homogeneity achieved on Ember was 82.77%, which is still a very decent result. On all of the datasets, V-Measure stayed relatively low (8% to 16%) for the best results - a consequence of penalization through Completeness because of the utilization of more clusters than the reported number of malware families.
@@ 4.2 Ablation Experiment on the Number of Components
The results of the ablation experiment are reported in Table 4. For Bodmas, the Homogeneity seemed to increase with every increase in the number of components except for the case of Autoencoder and 30 components. For Ember, Homogeneity (almost) consistently increased with the increase in the number of components, with the exception of “H-train-30” where it slightly dropped. The results for the Security dataset are the least consistent among the three datasets. In the case of PCA and 30 components, the Homogeneity decreased a little. On the other hand, for Autoencoder and 30 components, we saw an increase of more than 2% Homogeneity. For the case of 30 components, the Homogeneity increased in four out of six cases (of combinations of dataset and representation). The Homogeneity increased in all six cases for 50 components (compared to the 10-component baseline). Conclusions about this experiment will be drawn in Section 5.
@@ 5.1 Interpretation of Results
The results presented in Section 4 provide multiple insights. The results clearly show that Bodmas was the easiest dataset to cluster, achieving above 92% Homogeneity for the top result. Contrastingly, Ember achieved significantly worse results in terms of Homogeneity (i.e., only reaching around 82% for the best result on Ember). This constitutes a considerable variance in results. Both of these datasets are public benchmark datasets and share the same set of $2381$ Ember features. Moreover, they were gathered only roughly 2 years apart - from January 2017 to March 2018 for Ember and from August 2019 to September 2020 for Bodmas [yang2021bodmas]. Therefore, the main differentiating factor are the samples inside the datasets. Thus, our key finding is that malware clustering quality is highly dependent on the composition of the dataset.
Our work modified the malware clustering scenario by including benign samples. We can make conclusions based on the results observed in our experiments and in the experiments found in the existing work presented in Section 2.Although often performed on much smaller datasets with much fewer malware families, the existing work showed 98% Purity, Precision between 72% to 89%, 92% V-Measure, and Homogeneity of 73% and 79%. Thus, our observed results for Bodmas, Ember and Security align with the existing literature’s observations. We conclude that the inclusion of benign samples did not have a dramatic effect on the clustering quality in terms of Homogeneity.
@@
Third, one of our primary motivations behind the experiment was also to see the impact of the increase in the number of components on the Ember dataset. Even for Ember, for which we have seen worse results in terms Homogeneity than Bodmas (in the main experiment), the increase in the number of components did not help much in increasing Homogeneity. Ideally, we would like to see that the number of components has a greater influence (i.e., at least 5%, ideally at least 10%) on the result Homogeneity, but this was not observed. This strengthens our previously stated argument that the influence of the samples (picked into the dataset) on the result Homogeneity cannot be overstated.
@@ 5.3 Future Work
A deeper investigation into the problem of worse results on the Ember dataset than on Bodmas is equally important. For instance, merging benign and malware samples from different datasets (e.g., Ember and Bodmas) and repeating the clustering experiments could offer better insights into malware clustering. This approach may help clarify the challenges associated with clustering different parts of these datasets.
@@ 6 Conclusion
We have established the state-of-the-art in malware clustering on large public malware benchmark datasets. The best results for Bodmas and Ember were achieved using K-Means and were 92% and 82% Homogeneity, respectively. Our results also indicate that K-Means and BIRCH achieve higher Homogeneity on these large datasets than DBSCAN and Hierarchical Agglomerative Clustering. Moreover, our findings show that the quality of malware clustering can vary considerably, even if they share the same set of features. This led us to conclude that malware clustering quality is dependent on the composition of the dataset. To examine the results more closely, we performed an ablation study, which increased the number of components of PCA and Autoencoder representations. The ablation results show that the increase in the number of components does not have a substantial effect on clustering Homogeneity, strengthening the result regarding dataset composition.